A torque of 4Nm is applied to accelerate a wheel from rest to 6rev/sec in a time of 10sec. Determine
(i) angular acceleration
(ii) angular velocity
(iii) angular momentum
(iv) angular impulse
(v) power applied
1
Expert's answer
2020-06-24T11:21:04-0400

Solution.

"M=4Nm;"

"\omega_0=0rev/s;"

"\omega=6rev/s;"

"t=10s;"

"\epsilon=\dfrac{\omega-\omega_0}{t};"

"\epsilon=\dfrac{6rev/s}{10}=0.6rew/s^2=3.768rad/s^2;"

"\omega=6rev/s=37.68rad/s;"

"M=I\epsilon\implies I=\dfrac{M}{\epsilon};"

"I=\dfrac{4Nm}{3.768rad/s^2}=1.06kgm^2;"

"L=I\omega;"

"L=1.06kgm^2\sdot37.68rad/s=40kgm^2/s;"

"P=M\omega;"

"P=4Nm\sdot37.68rad/s=150.72W" ;

Answer: "i)\epsilon=3.768rad/s;"

"ii)\omega=37.68rad/s;"

"iii)I=1.06kgm^2;"

"iv)L=40kgm^2\/s;"

"v)P=150.72W;"